# Ultrafast Sampling-based Kinodynamic Planning via Differential Flatness

### 17 Mar 2026 v1

@@ Abstract
Motion planning under dynamics constraints, i.e., kinodynamic planning, enables safe robot operation by generating dynamically feasible trajectories that the robot can accurately track. For high-d o f robots such as manipulators, sampling-based motion planners are commonly used, especially for complex tasks in cluttered environments. However, enforcing constraints on robot dynamics in such planners requires solving either challenging two-point boundary value problems (BVPs) or propagating robot dynamics over time, both of which are computational bottlenecks that drastically increase planning times. Meanwhile, recent efforts have shown that sampling-based motion planners can generate plans in microseconds using parallelization, but are limited to geometric paths. This paper develops AkinoPDF, a fast parallelized sampling-based kinodynamic motion planning technique for a broad class of differentially flat robot systems, including manipulators, ground and aerial vehicles, and more. Differential flatness allows us to transform the motion planning problem from the original state space to a flat output space, where an analytical time-parameterized solution of the BVP and dynamics integration can be obtained. A trajectory in the flat output space is then converted back to a closed-form dynamically feasible trajectory in the original state space, enabling fast validation via “single instruction, multiple data” parallelism. Our method is fast, exact, and compatible with any sampling-based motion planner. We extensively verify the effectiveness of our approach in both simulated benchmarks and real experiments with cluttered and dynamic environments, requiring mere microseconds to milliseconds of planning time.
@@ I Introduction
Motion planning is critical for safe and accurate robot operation in many applications, e.g., transportation [claussmann2020review], environment monitoring [honig2018trajectory, zhou2022swarm], warehouses [eppner2016lessons], healthcare [riek2017healthcare], and home assistance [jenamani2025feast]. In such applications, the robot has to react quickly to changes in its environment, promptly (re)plan a collision-free trajectory, and safely track the trajectory, using a controller, to reach a goal state. This task requires fast motion planning, subject to both collision avoidance and robot dynamics constraints, to generate a dynamically feasible trajectory from a start to a goal that the robot is able to follow. While recent advances have significantly reduced motion planning times for geometric planning via parallelization techniques [thomason2024vamp, ramsey2024-capt], sampling-based motion planning under dynamics constraints, i.e., kinodynamic planning, remains a challenge for real-time applications especially for high degree-of-freedom (d o f) robots such as manipulators. In this paper, we address this problem by leveraging the differential flatness property of many common robot systems, such as ground and aerial vehicles, manipulators, and more, to enable ultrafast sampling-based kinodynamic motion planning via parallelization.
@@
Figure 1: Motion planning for a “pick and place” task in a cluttered environment with narrow passage: a dynamically feasible trajectory (a) generated from our kinodynamic planner (AkinoPDF) can be accurately tracked by a UR5 robot. Meanwhile, tracking a geometric path (b) leads to collisions (shown in red) that topple the nearby boxes. Multiple intermediate states are overlaid to illustrate the robot’s motion. Our planner is real-time and generates trajectories in $\sim 90\mu s$ by leveraging differential flatness and “single instruction, multiple data” (simd) parallelism.
A key challenge in enforcing dynamics constraints in sampling-based planning is to solve challenging two-point boundary value problems (BVPs) for dynamically feasible local paths between robot states. Many planners [lavalle2001randomized, hsu2022randomizekinodynamic, li2016sst] avoid this by instead sampling a control input and then propagating the dynamics forward using numerical integration [butcher2016numerical]. Other works solve the BVP problem for simple systems [karaman2010kinorrt] and linearized robot dynamics [webb2013kinodynamicRRT, Perez2012LQRRRT], or approximate the solution using a neural network [wolfslag2018rrt, chiang2019rl, zheng2021sampling]. Instead, we obtain an exact analytical solution of the BVP and dynamics propagation, thanks to the differential flatness of many robot platforms such as mobile robots and manipulators [murray1995differential, mellinger2011minimumsnap].
Differential flatness [murray1995differential] is a powerful system property that, similar to feedback linearization [khalil2002nonlinear], simplifies motion planning and control designs by converting the nonlinear robot dynamics to an equivalent linear system. For differentially flat systems, the robot states and control inputs can be described by a set of carefully chosen flat outputs and their derivatives (see Def. 1 for more details). A trajectory in the flat output space can be converted to a dynamically feasible trajectory in the robot’s original state space, allowing us to simplify motion planning problems via change of variables. This technique has been used mainly to plan trajectories for mobile robots, most notably quadrotors [mellinger2011minimumsnap, mellinger2012_trajgen, liu2017search]. However, the planning time is still hindered by computationally expensive subroutines such as forward kinematics and collision checking, which are particularly worse with high-d o f robots such as manipulators.
@@
In summary, we develop AkinoPDF, an “Accelerated Sampling-based Kinodynamic Planning via Differential Flatness” technique that:
@@
  • and performs fast collision checking of the closed-form trajectories using fine-grained simd parallelism.
We extensively verify our approach with challenging motion planning problems on low- and high-d o f, fully- and under-actuated robot systems in both simulated and real experiments.
@@ II-A Motion Planning with Dynamics Constraints
BVPs and dynamics propagation are both major computational bottlenecks for kinodynamic planning. While BVPs can be analytically solved for linear or simple systems [webb2013kinodynamicRRT, Perez2012LQRRRT], this is not true in general for most nonlinear systems and it is computationally expensive to obtain an approximate solution. Kinodynamic RRT [webb2013kinodynamicRRT, Perez2012LQRRRT] linearizes the dynamics around an operating point and demonstrates that BVPs can be solved in closed form for certain robots, e.g., quadrotors around a hovering position. However, the approximated solution only works well around the operating point, limiting aggressive maneuvers. Meanwhile, dynamics propagation, given a sampled control input, often requires numerical integration such as Euler’s or Runge-Kutta methods [butcher2016numerical] to sequentially calculate the next robot state over multiple small time steps, which is time consuming and inimical to parallelism.
@@
Optimization-based, search-based and sampling-based approaches can be combined to improve trajectory generation [ortizharo2024iDbRRT, ortizharo2025iDbAstar, natarajan2024pinsat, natarajan2023torque, natarajan2021interleaving, sakcak2019sampling, shome2021asymptotically, kamat2022bitkomo, choudhury2016regionally, alwala2021joint]. For example, a path or trajectory from sampling-based or search-based planners can be used as an initial solution for optimization-based ones [ortizharo2024iDbRRT, ortizharo2025iDbAstar]. INSAT planners [natarajan2024pinsat, natarajan2023torque, natarajan2021interleaving] interleave between search-based planning on a low dimensional subspace and optimization-based planning on the full-dimensional space to improve planning times and success rates. Meanwhile, a library of precomputed motion primitives from search-based planning can be sampled to expand the planning tree in sampling-based motion planning [sakcak2019sampling, shome2021asymptotically]. Furthermore, optimized local paths can be used to generate collision-free edges and bias the sampling regions in a sampling-based planner [kamat2022bitkomo, choudhury2016regionally]. Another approach is to fit a geometric path with a time-parameterized trajectory using trajectory optimization such as TOPP-RA [pham2018toppra] or Ruckig [berscheid2021ruckig], typically without considering collision avoidance constraints.
@@
Our method performs fast sampling-based kinodynamic planning by leveraging simd parallelism (Sec. II-B) and the differential flatness of the dynamics of common robot platforms [allen2019real, liu2017search, bascetta2017flat, welde2021dynamically] to directly tackle the BVP and dynamics propagation problems. A system is called differentially flat if there exist variables, called flat outputs, whose values and derivatives dictate the robot state and control inputs (see Def. 1 for details). This approach has been applied to sampling-based [bascetta2017flat, ye2022efficient, han2023efficient], search-based [liu2017search], and optimization-based motion planning [mellinger2011minimumsnap, welde2021dynamically, han2023efficient, hao2005differential], but primarily for low-dimensional robot platforms, e.g., quadrotors, or for a specific system, e.g., gantry cranes [vu2022sampling]. Instead, we integrate differential flatness with sampling-based kinodynamic planning for generic high-d o f robots, where forward kinematics and collision checking are time-consuming besides enforcing dynamics constraints. While most existing methods approximate a solution to the BVP problem and dynamics propagation, differential flatness allows us to obtain a closed-form fixed- or minimum-time polynomial solution. Such closed-form trajectories are amenable to parallelized collision checking using fine-grained parallelization based on simd instructions (Sec. II-B), enabling trajectory generation in microseconds to milliseconds.
@@ II-B Hardware-accelerated Motion Planning
Fine-grained parallelization techniques focus on parallelizing primitive operations in a motion planning algorithm, e.g., via “single instruction/multiple data” (simd) instructions on consumer-grade CPUs. This has been shown to provide extremely fast geometric motion planning subject to collision avoidance constraints, with planning times ranging from microseconds to milliseconds [thomason2024vamp, ramsey2024-capt, wilson2024nearest]. To check an edge (a line segment) for collision, VAMP [thomason2024vamp] uses simd instructions to efficiently perform parallelized forward kinematics and collision checking on multiple configurations, generated via linear interpolation. While this approach is promising, it is challenging to enforce additional requirements via parallelized primitive operations such as dynamics and non-holonomic constraints. As mentioned in Sec. II-A, constructing a robot dynamically feasible trajectory by solving a BVP or dynamics propagation requires either approximation or sequential numerical approximation, which is, by design, not compatible with fine-grained parallelism. We instead leverage the differential flatness property to obtain time-parameterized solutions that can be discretized at arbitrary times and hence, amenable to SIMD-based primitive operations.
@@ III Problem Formulation
$\dot{\boldsymbol{\sigma}}=\mathbf{f}(\boldsymbol{\sigma},\mathbf{u}).$ (2)
Given an initial robot state $\mathbf{x}_{s}$ and a goal region ${\cal G}\in{\cal X}$, the kinodynamic motion planning problem aims to find a dynamically feasible trajectory $\boldsymbol{\sigma}(t)$ with control input $\mathbf{u}(t)$, connecting the initial state $\mathbf{x}_{s}$ to the goal region $\cal G$} in the free space ${\cal X}_{free}$ as described in Problem 1.
@@
Given an initial robot state $\mathbf{x}_{s}$ and a goal region $\cal G$}, find a control input $\mathbf{u}(t)$ that generates a dynamically feasible trajectory $\boldsymbol{\sigma}(t)$ such that:
$\dot{\boldsymbol{\sigma}}=\mathbf{f}(\boldsymbol{\sigma},\mathbf{u}),$ (3)
@@
In the remainder of the paper, we solve Problem 1 by developing a parallelized kinodynamic motion planning technique for differential flat robot systems, including common platforms such as ground and aerial vehicles, and manipulators. Our approach offers ultrafast planning time and exact dynamically feasible trajectory solution without the need to approximate the robot dynamics. Occasionally, we will drop the notation of time dependence for readability.
@@ IV-A Sampling-based Kinodynamic Motion Planning
Propagation-based expansion of the search graph, however, severely limits the performance of planners, as the local path $\mathbf{x}_{loc}(t)$ is often approximated by Euler’s or Runge-Kutta’s integration methods over a sequence of small time steps for high accuracy.
@@ IV-C CPU fine-grained parallelism
Fine-grained parallelization techniques such as “single instruction, multiple data” (simd) are ubiquitous on consumer CPUs and have recently shown significant improvement in sampling-based geometric motion planning by performing forward kinematics and collision checking on multiple configurations in parallel. Vectorization via simd allows simultaneously applying the same primitive operations on multiple variables. VAMP [thomason2024vamp], a simd-accelerated planning method, uses this technique to perform parallelized collision checking of multiple robot configurations, evenly sampled along a line segment (Fig. 2(a)) connecting two nodes on a planning tree or graph. Given the set of configurations, the robot’s geometric shape, described by a set of spheres in the workspace, is calculated via branchless forward kinematics and checked for collision against the obstacle geometry.
simd-based collision checking requires access to all the states along a local path, typical through an analytical form of the motion. However, under dynamics constraints, the nonlinear local path $\mathbf{x}_{loc}(t)$ that solves the BVP problem or dynamics propagation (Sec. IV-A) is either computationally intractable or can only be sequentially approximated, prohibiting configuration sampling at multiple arbitrary times in advance. In the next section, we address this problem by deriving a closed-form time-parameterized local path $\mathbf{x}_{loc}(t)$ based on the differential flatness property and sampling multiple states at different times, as illustrated in Fig. 2(b).

### 21 Apr 2026 v2

@@ Abstract
Motion planning under dynamics constraints, i.e., kinodynamic planning, enables safe robot operation by generating dynamically feasible trajectories that the robot can accurately track. For high-d o f robots such as manipulators, sampling-based motion planners are commonly used, especially for complex tasks in cluttered environments. However, enforcing constraints on robot dynamics in such planners requires solving either challenging two-point boundary value problems (BVPs) or propagating robot dynamics, both of which cause computational bottlenecks that drastically increase planning times. Meanwhile, recent efforts have shown that sampling-based motion planners can generate plans in microseconds using parallelization, but are limited to geometric paths. This paper develops FLASK, a fast parallelized sampling-based kinodynamic motion planning framework for a broad class of differentially flat robot systems, including manipulators, ground and aerial vehicles, and more. Differential flatness allows us to transform the motion planning problem from the original state space to a flat output space, where an analytical time-parameterized solution of the BVP problem can be obtained. A trajectory in the flat output space is then converted back to a closed-form dynamically feasible trajectory in the original state space, enabling fast validation via “single instruction, multiple data” parallelism. Our framework is fast, exact, and compatible with any sampling-based motion planner, while offering theoretical guarantees on probabilistic exhaustibility and asymptotic optimality based on the closed-form BVP solutions. We extensively verify the effectiveness of our approach in both simulated benchmarks and real experiments with cluttered and dynamic environments, requiring mere microseconds to milliseconds of planning time.
@@ I Introduction
Motion planning is critical for safe and accurate robot operation in many applications, e.g., transportation [claussmann2020review], environment monitoring [honig2018trajectory, zhou2022swarm], warehouses [eppner2016lessons], healthcare [riek2017healthcare], and home assistance [jenamani2025feast]. In such applications, the robot has to react quickly to changes in its environment, promptly (re)plan a collision-free trajectory, and safely track the trajectory to reach a goal state, using a controller. This task requires fast motion planning, subject to both collision avoidance and robot dynamics constraints, to generate a dynamically feasible trajectory from a start to a goal that the robot is able to follow. While recent advances have significantly reduced motion planning times for geometric planning via parallelization techniques [thomason2024vamp, ramsey2024-capt], sampling-based motion planning under dynamics constraints, i.e., kinodynamic planning, remains a challenge for real-time applications especially for high degree-of-freedom (d o f) robots such as manipulators. In this paper, we address this problem by leveraging the differential flatness property of many common robot systems, such as ground and aerial vehicles, manipulators, and more, to enable ultrafast sampling-based kinodynamic motion planning via parallelization techniques.
@@
Figure 1: Motion planning for a “pick and place” task in a cluttered environment: a dynamically feasible trajectory (a) generated from our FLASK framework can be accurately tracked by a UR5 robot. Meanwhile, tracking a geometric path (b) leads to collisions (shown in red) that topple the nearby boxes. Multiple intermediate states are overlaid to illustrate the robot’s motion. Our planning framework is real-time and generates trajectories in $\sim 90\mu s$ by leveraging differential flatness and “single instruction, multiple data” (simd) parallelism.
A key challenge in enforcing dynamics constraints in sampling-based planning is to solve challenging two-point boundary value problems (BVPs) for dynamically feasible local paths between robot states. Many planners [lavalle2001randomized, hsu2022randomizekinodynamic, li2016sst] avoid this by instead sampling an often sub-optimal control input and then propagating the dynamics forward using numerical integration [butcher2016numerical]. Dynamic propagation with randomly sampled control often causes the planner to “wander” to irrelevant parts of the state space and might lead to longer paths and longer planning times. Other works solve the BVP problem for simple systems [karaman2010kinorrt] and linearized robot dynamics [webb2013kinodynamicRRT, Perez2012LQRRRT], or approximate the solution using a neural network [wolfslag2018rrt, chiang2019rl, zheng2021sampling]. Instead, we obtain an exact analytical solution of the BVP, thanks to the differential flatness of many robot platforms such as mobile robots and manipulators [murray1995differential, mellinger2011minimumsnap].
Differential flatness [murray1995differential] is a powerful system property that, similar to feedback linearization [khalil2002nonlinear], simplifies motion planning and control designs by converting the nonlinear robot dynamics to an equivalent linear system. For differentially flat systems, the robot states and control inputs can be described by a set of carefully chosen flat outputs and their derivatives (see Def. 1 for more details). A trajectory in the flat output space can be converted to a dynamically feasible trajectory in the robot’s original state space, allowing us to simplify motion planning problems via a change of variables. This technique has been used mainly to plan trajectories for mobile robots, most notably quadrotors [mellinger2011minimumsnap, mellinger2012_trajgen, liu2017search]. However, the planning time is still hindered by computationally expensive subroutines such as forward kinematics and collision checking, which are particularly worse with high-d o f robots such as manipulators.
@@
In summary, we develop FLASK  [^1] The code will be made publicly available., a “Differential Flatness-based Accelerated Sampling-based Kinodynamic Planning” framework that:
@@
  • performs fast collision checking of the closed-form trajectories using fine-grained simd parallelism.
We provide a theoretical analysis on our approach’s probabilistic exhausitivity, a concept stronger than completeness introduced in [schmerling2015optimal_driftless, kavraki1998analysis], and optimality guarantees. We extensively verify our approach with challenging motion planning problems on low- and high-d o f, fully- and under-actuated robot systems in both simulated and real experiments. FLASK is shown to achieve planning times of merely a few milliseconds, even on high-d o f robot platforms in cluttered environments, while respecting dynamics constraints by design. Our framework is general and can be integrated in the context of many common sampling-based motion planners, such as RRT-Connect [kuffner2000rrtconnect], SST* [li2016sst] and others, effectively turning geometric planners into kinodynamic planners with theoretical guarantees.
@@ II-A Motion Planning with Dynamics Constraints
BVPs and dynamics propagation cause major computational bottlenecks for kinodynamic planning. While BVPs can be analytically solved for linear or simple systems [webb2013kinodynamicRRT, Perez2012LQRRRT], this is not true in general for most nonlinear systems and it is computationally expensive to obtain an approximate solution. Kinodynamic RRT [webb2013kinodynamicRRT, Perez2012LQRRRT] linearizes the dynamics around an operating point and demonstrates that BVPs can be solved in closed form for certain robots, e.g., quadrotors around a hovering position. However, the approximated solution only works well around the operating point, limiting aggressive maneuvers. To avoid solving difficult BVPs, most existing kinodynamic planners [li2016sst] resort to dynamics propagation, where a constant value of the control is sampled instead. A numerical integrator such as Euler’s or Runge-Kutta methods [butcher2016numerical] is then used to sequentially calculate the next robot state over multiple small time steps to maintain high accuracy. Typically, the sampled control is suboptimal, and therefore, causes the planner to wander in the state space before reaching the goal. Existing planners often mitigate this “wandering” effect via best-state selection and tree pruning [li2016sst], but still require long planning times to find a dynamically feasible trajectory.
@@
Optimization-based, search-based and sampling-based approaches can be combined to improve trajectory generation [ortizharo2024iDbRRT, ortizharo2025iDbAstar, natarajan2024pinsat, natarajan2023torque, natarajan2021interleaving, sakcak2019sampling, shome2021asymptotically, kamat2022bitkomo, choudhury2016regionally, alwala2021joint]. For example, a path or trajectory from sampling-based or search-based planners can be used as an initial solution for optimization-based ones [ortizharo2024iDbRRT, ortizharo2025iDbAstar]. INSAT planners [natarajan2024pinsat, natarajan2023torque, natarajan2021interleaving] interleave between search-based planning on a low-dimensional subspace and optimization-based planning on the full-dimensional space to improve planning times and success rates. Meanwhile, a library of precomputed motion primitives from search-based planning can be sampled to expand the planning tree in sampling-based motion planning [sakcak2019sampling, shome2021asymptotically]. Furthermore, optimized local paths can be used to generate collision-free edges and bias the sampling regions in a sampling-based planner [kamat2022bitkomo, choudhury2016regionally]. Another approach is to fit a geometric path with a time-parameterized trajectory using trajectory optimization such as TOPP-RA [pham2018toppra] or Ruckig [berscheid2021ruckig], typically without considering collision avoidance constraints.
@@
Our method performs fast sampling-based kinodynamic planning by leveraging simd parallelism (Sec. II-B) and the differential flatness of the dynamics of common robot platforms [allen2019real, liu2017search, bascetta2017flat, welde2021dynamically] to directly tackle the BVP problems. A system is called differentially flat if there exist variables, called flat outputs, whose values and derivatives dictate the robot state and control inputs (see Def. 1 for details). This approach has been applied to sampling-based [bascetta2017flat, ye2022efficient, wang2024differential, seemann2014exact], search-based [liu2017search], and optimization-based motion planning [mellinger2011minimumsnap, welde2021dynamically, han2023efficient, hao2005differential, beaver2024optimal], but primarily for low-dimensional robot platforms (often with simple geometry shapes such as spheres or boxes for collision checking), e.g., quadrotors [mellinger2011minimumsnap], unicycles and 2-link arms [beaver2024optimal], or for a specific system, e.g., gantry cranes [vu2022sampling]. Instead, we integrate differential flatness with sampling-based kinodynamic planning for generic high-d o f robots, where forward kinematics and collision checking are complex and time-consuming besides enforcing dynamics constraints. While most existing methods approximate a solution to the BVP problems, differential flatness allows us to obtain a closed-form fixed- or minimum-time polynomial BVP solution. Such closed-form trajectories are amenable to parallelized collision checking using fine-grained parallelization based on simd instructions (Sec. II-B), enabling trajectory generation in microseconds to milliseconds. It is also important to note that existing flatness-based sampling-based planners either do not prove completeness and optimality or loosely mention the existing guarantees of geometric planners, e.g., those of RRT* [bascetta2017flat, vu2022sampling, wang2024differential, seemann2014exact, ye2022efficient]. Achieving such theoretical guarantees with closed-form BVP solutions is nontrivial as it requires careful consideration of nonlinear local paths rather than linear edges on the graph. We address this by offering a theoretical analysis of probabilistic exhaustivity, which is stronger than completeness, and asymptotic optimality of our approach in Sec. VI.
@@ II-B Hardware-accelerated Motion Planning
Fine-grained parallelization techniques focus on parallelizing primitive operations in a motion planning algorithm, e.g., via “single instruction/multiple data” (simd) instructions on consumer-grade CPUs. This has been shown to provide extremely fast geometric motion planning subject to collision avoidance constraints, with planning times ranging from microseconds to milliseconds [thomason2024vamp, ramsey2024-capt, wilson2024nearest]. To check an edge (a line segment) for collision, VAMP [thomason2024vamp] uses simd instructions to efficiently perform parallelized forward kinematics and collision checking on multiple configurations, generated via linear interpolation. While this approach is promising, it is challenging to enforce additional requirements via parallelized primitive operations such as dynamics and non-holonomic constraints. Particularly, enforcing nonlinear dynamics constraints with fine-grained parallelism is non-trivial due to the intractability of BVP problems, as mentioned in Sec. II-A. We instead leverage the differential flatness property to obtain time-parameterized solutions that can be discretized at arbitrary times and hence, amenable to SIMD-based primitive operations. Our deliberate fusion of differential flatness and SIMD parallelism effectively brings the benefits of “fine-grained” parallelized forward kinematics and collision checking to kinodynamic planning. Thanks to the closed-form BVP solution in the flat output space, we will later show that our approach achieves planning times of merely a few milliseconds.
@@ III Problem Formulation
$\dot{\boldsymbol{\sigma}}{(t)}=\mathbf{f}(\boldsymbol{\sigma}{(t)},\mathbf{u}{(t)}),\quad{\forall t\in[0,1]}.$ (2)
Given an initial robot state $\mathbf{x}_{s}$ and a goal region ${\cal G}{\subset{\cal X}}$, as a subset of $\cal X$}, the kinodynamic motion planning problem aims to find a dynamically feasible trajectory $\boldsymbol{\sigma}(t)$ with control input $\mathbf{u}(t)$, connecting the initial state $\mathbf{x}_{s}$ to the goal region $\cal G$} in the free space ${\cal X}_{free}$ as described in Problem 1. The goal region $\cal G$} commonly represents a desired state or a region that a goal state can be sampled from.
@@
Given an initial robot state $\mathbf{x}_{s}$ and a goal region ${\cal G}{\subset{\cal X}}$, find a control input $\mathbf{u}(t)$ that generates a dynamically feasible trajectory $\boldsymbol{\sigma}(t)$ such that:
$\dot{\boldsymbol{\sigma}}{(t)}=\mathbf{f}(\boldsymbol{\sigma}{(t)},\mathbf{u}{(t)}),$ (3)
@@
In the remainder of the paper, we solve Problem 1 by developing FLASK, a parallelized kinodynamic motion planning framework for differential flat robot systems, including common platforms such as ground and aerial vehicles, and manipulators. Our approach offers ultrafast planning time and exact dynamically feasible trajectory solution without the need to approximate the robot dynamics. Occasionally, we will drop the notation of time dependence for readability.
@@ IV-C CPU fine-grained parallelism
Fine-grained parallelization techniques such as “single instruction, multiple data” (simd) are ubiquitous on consumer CPUs and have recently shown significant improvement in sampling-based geometric motion planning by performing forward kinematics and collision checking on multiple configurations in parallel. Vectorization via simd allows simultaneously applying the same primitive operations on multiple variables. VAMP [thomason2024vamp], a simd-accelerated planning method, uses this technique to perform parallelized collision checking of multiple robot configurations, evenly sampled along a line segment (Fig. 2(a)) connecting two nodes on a planning tree or graph. Given the set of configurations, the robot’s geometric shape, described by a set of SIMD-compatible shapes such as capsules or spheres in the workspace, is calculated via branchless forward kinematics and checked for collision against the obstacle geometry.
simd-based collision checking requires access to all the states along a local path, typically through an analytical form of the motion. However, under dynamics constraints, the nonlinear local path $\mathbf{x}_{loc}(t)$ that solves the BVP problem (Sec. IV-A) is computationally intractable, prohibiting configuration sampling at multiple arbitrary times in advance. In the next section, we address this problem by deriving a closed-form time-parameterized local path $\mathbf{x}_{loc}(t)$ based on the differential flatness property and sampling multiple states at different times, as illustrated in Fig. 2(b).